Mark is twice as old as Dave. In four years the sum of their ages will be 26. How old is Mark today?
Answer by JBarnum(2146)

Mark is twice as old as Dave. In four years the sum of their ages will be 26. How old is Mark today?
M=2D
(M+4)+(D+4)=26, M+D+8=26, M+D=18
2D+D=18
3D=18
D=6
M=2(6)
M=12
M=Mark
D=Dave